OpenStudy (anonymous):

Is the cotan(theta) function the same as tan^-1(theta)?

OpenStudy (anonymous):

\[\cot \theta = \frac{ 1 }{ \tan \theta }\]

OpenStudy (anonymous):

So that's a yes? If I'm entering it into my calculator (there is no cot function), I enter\[\cot(\theta) = \tan^{-1}(\theta)\]

OpenStudy (anonymous):

you can't write it as \[\tan ^{-1} \theta \]

OpenStudy (anonymous):

Isn't that the same thing...? Doesn't: \[\frac{ 1 }{ \tan(\theta) } = \tan ^{-1}(\theta)\]

OpenStudy (anonymous):

no my friend

OpenStudy (anonymous):

:( okay. Thanks

OpenStudy (anonymous):

nope. to find cot(theta), do this: 1/(tan(theta))

OpenStudy (anonymous):

when we take tan function from one side of equality to other then we write tan^-1

OpenStudy (anonymous):

So you use tan^-1 in something like... To solve for x if tan(x)=a, then x=tan^-1(a)

OpenStudy (anonymous):

yes you are right

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

Can you say\[\csc ^{4}(\theta) = \frac{ 1 }{ \sin ^{4}(\theta) }\]

OpenStudy (anonymous):

yep. in calculator, that would be: 1/(sin(theta))^4

OpenStudy (shubhamsrg):

-1 is a special case in trigonometric functions.. -1 corresponds to inverse function.. thus for 1/sin, we use cosec and likewise..

OpenStudy (anonymous):

Thanks I get it now :)

OpenStudy (shubhamsrg):

hmm,,glad you do!